Avec Úvo SRE L VII. (5pt) Consider the above-right common-collector or emitter-follower BJT amplifier circuit. Given: ß= 100, RE = 10 k1, Vcc
Answer #1

D Caui valeot Circut: Vec 20 1야 BYKL in inpdttoop 20 t (T t)R, RB te)eEo 20 (T+)R+ o7 +( lox 0 EOr = 113 ( o.7 =alyDA = 0-7 .Small sin Escivalen FF EL VD Smll sta mocde RL FE MCA) CitpRe (101)2K11e) Is46K +Go1)RK1OK) Cro1)1.66k) Sy6x +0166x) 6 33 : ০৭৭ r69. 79 K Av AvP ITAV,B RI tok PitrB Vo RitRe 1o(e) Cument gn RDut V.C Zin Av out 284 169. 264 Av AS Power gan O 576x169.26y